OpenStudy (anonymous):

Solve for a 24a+48=18a-36

OpenStudy (anonymous):

Try not to have funny looking numbers on separate sides. They like to be together. 24a + 48 = 18a - 36. 24 is larger than 18, so I'd prefer to move the 18 to the left side. 24a - 18a + 48 = -36. 6a + 48 = -36. Unfortunately, we're going to have to subtract the 48, so we'll need to do -36 - 48, which is the same thing as -36 + -48, and that equals -84. 6a = -84 a=-84/6 a=-14